OpenStudy (maggy14145):

Given p(x) = 2x3 − 8x2 + 13x − 21, ___ is a factor because ___.

OpenStudy (maggy14145):

The options are: A. p(3) does not equal zero B. p(-3) does not equal zero C. x-3 D. p(3)=0 E. x+3 F. p(-3) = 0 And I'm supposed to pick two from the above to fill in the blanks!

OpenStudy (maggy14145):

For the first blank I believe the answer should be C. because x-3 is one of the factors from the equation but I'm confused about the second blank

OpenStudy (maggy14145):

SOMEONE PLEASE HELP

OpenStudy (prathamesh_m):

Remember this : If x-a is a factor of p(x) then p(a)=0.

OpenStudy (prathamesh_m):

Can you guess why?

OpenStudy (maggy14145):

Ss then the answer for the second one would be p(3)=0 ?

OpenStudy (maggy14145):

since technically you're supposed to set the factor equal to zero..?

OpenStudy (prathamesh_m):

Yes that's right

OpenStudy (maggy14145):

thank you so much!

OpenStudy (prathamesh_m):

You're welcome!!

OpenStudy (maggy14145):

can u help me with a graphing problem as well?

OpenStudy (maggy14145):

Gerry wrote a new function, g(x), based on the parent function f(x) = √x. To create g(x), she replaced x with x + k, so that g(x) = f(x+ k) and got the graph shown below. Based on the graph, k = _____.

OpenStudy (maggy14145):

http://cds.flipswitch.com/tools/asset/media/384951

OpenStudy (welshfella):

this transformation is a translation of the whole graph k units horizontally to the left. note that f(x) will pas through the origin (0,0)

OpenStudy (welshfella):

so how many units has f(x) been translated ?

OpenStudy (maggy14145):

7 units to the left I believe..?

OpenStudy (welshfella):

correct so k = 7

OpenStudy (maggy14145):

oh..was it really that easy?

OpenStudy (welshfella):

yes

OpenStudy (maggy14145):

so the final answer is k=7?

OpenStudy (welshfella):

right

OpenStudy (maggy14145):

wow. okay. I feel stupid lol. thank you so much for justifying this for me. I greatly appreciate it :)

OpenStudy (welshfella):

https://www.desmos.com/calculator/gdfnetsubu - there are the 2 graphs

OpenStudy (maggy14145):

could you help me with another question please?

OpenStudy (maggy14145):

Rachel knows that the value of b in a quadratic function can influence the axis of symmetry as well as the vertex of the function. She wants to know how it affects the x-intercepts of the function as well, so she starts by substituting y = 0 into the standard form for a quadratic function. She tries to rewrite the quadratic equation to solve for b. Her work, and answers, are shown below. What mistake, if any, did she make?

OpenStudy (maggy14145):

http://cds.flipswitch.com/tools/asset/media/384717

OpenStudy (welshfella):

can you see why step 2 is wrong? at this stage she needs to isolate bx

OpenStudy (welshfella):

i have to go foe about 10 minutes

OpenStudy (maggy14145):

ahh rachel should have subtracted c instead of adding it